One leg of a right triangle is 1 foot longer than the other leg. The hypotenuse is 9 feet longer than the shorter leg. Find the length of the sides of the triangle.
Answer by mananth(16946)

the three sides of the right triangle are
x, x+1, x+9
x^2+(x+1)^2=(x+9)^2
x^2+x^2+2x+1 = x^2+18x+81
x^2-16x-80=0
x^2-20x+4x-80=0
x(x-20)+4(x-20)=0
(x-20)(x+4)=0
x=20 which is positive
20 21,& 29 are the sides
m.ananth@hotmail.ca